# Characterizing Gaussian quantum processes with Gaussian resources

### v1

@@ VIII Conclusion
We have presented a method for characterizing Gaussian quantum processes using only Gaussian resources. Our approach directly reconstructs the symplectic matrix that describes an arbitrary Gaussian process, requiring only $O(N^{2})$ measurements for an $N$-mode system. As this method uses only experimentally accessible resources such as coherent probes and quadrature detection methods, it is well-suited for implementation in current experimental setups. A highlight of this method is that it is resilient to uniform loss.
@@ Appendix A Unitary and symplectic descriptions of linear optics
Consider a $N-$mode linear optical device with $\hat{a}_{i}^{\dagger}\,(\hat{b}_{i}^{\dagger})$ as the input (output) modes, where $i\in\{1,2,\ldots,N\}$. Let $\bf{\hat{a}^{\dagger}},(\hat{b}^{\dagger})$ be the vector of the input (output) mode operators. The input and output quadrature operators are defined from the mode creation and annihilation operators as
@@ Appendix B Uniform loss model in Gaussian devices
Consider a Gaussian quantum device with uniform loss in all modes. This loss is modeled by a set of fictitious beam splitters with the same transmissivity $\eta$ in each mode. We demonstrate that this type of loss commutes with the device, meaning that all losses can be effectively moved to the beginning of the device. This was shown to be true in Ref. Oszmaniec and Brod (2018) for the restricted case of passive linear optics.

### v2

@@ VII.1 Superiority of heterodyne measurements
VII.2 Detecting non-Gaussian dynamics
Having shown that the proposed method can characterize Gaussian dynamics efficiently, we now show that it can also detect if the dynamics is beyond the Gaussian regime. To this end, we consider a simple model of the non-Gaussian dynamics in a single-mode cubic phase gate $e^{i\gamma\hat{X}^{3}}$ [32]. Such systems lead to the transformation of the quadrature operators nonlinearly as [33]
$\hat{X}^{\text{out}}$ $=\hat{X}^{\text{in}},$
$\hat{P}^{\text{out}}$ $=\hat{P}^{\text{in}}+3\gamma\left(\hat{X}^{\text{in}}\right)^{2}.$ (17)
This nonlinearity shall be apparent as we characterize the system by the proposed method, as the supposed symplectic matrix elements will be found to have elements that depend explicitly on the probe used. In this instance, when the system is probed with a real coherent state and the momentum quadrature is measured at the output, the ratio of the output quadrature to the input quadrature will explicitly depend on the input probe amplitude, $\it{viz.}$,
$\frac{\left\langle\hat{P}^{\text{out}}\right\rangle}{\left\langle\hat{X}^{\text{in}}\right\rangle}=3\gamma\left\langle\hat{X}^{\text{in}}\right\rangle.$ (18)
Therefore, repeating said process with a different coherent probe will indicate the presence nonlinearity beyond Gaussian dynamics.
@@ VIII Conclusion
We have presented a method for characterizing Gaussian quantum processes using only Gaussian resources. Our approach directly reconstructs the symplectic matrix that describes an arbitrary Gaussian process, requiring only $O(N^{2})$ measurements for an $N$-mode system. This method is scalable, and is resilient to optical loss and phase-modulation errors. As this method uses only experimentally accessible resources such as coherent probes and quadrature detection methods, it is well-suited for implementation in current experimental setups.
@@ Appendix A Unitary and symplectic descriptions of linear optics
Consider a NN-mode linear optical device with $\hat{a}_{i}^{\dagger}\,(\hat{b}_{i}^{\dagger})$ as the input (output) modes, where $i\in\{1,2,\ldots,N\}$. Let $\bf{\hat{a}^{\dagger}},(\hat{b}^{\dagger})$ be the vector of the input (output) mode operators. The input and output quadrature operators are defined from the mode creation and annihilation operators as
@@ Appendix B Uniform loss model in Gaussian devices
Consider a Gaussian quantum device with uniform loss in all modes. This loss is modeled by a set of fictitious beam splitters with the same transmissivity $\eta$ in each mode. We demonstrate that this type of loss commutes with the device, meaning that all losses can be effectively moved to the beginning of the device. This was shown to be true in Ref. [27] for the restricted case of passive linear optics.